# Coordination Control of Discrete Event Systems under Cyber Attacks

### v2

@@ 5 Coordination Control under Cyber Attacks without CA-Controllability and CA-Observability
Under Assumption 1, if $K$ is conditionally decomposable, then for any supervisors $\mathcal{S}_{i}$ for $G_{i}$, $i=1,2$, we have
@@
Step 4. Replace unobservable transitions in $H_{i}^{\uparrow,a}$ by $\varepsilon$-transitions to obtain
@@
Step 5. Construct the CA-observer of $H_{i,\varepsilon}^{\uparrow,a}$ by determinizing $H_{i,\varepsilon}^{\uparrow,a}$ as
Step 6. Design a CA-supervisor $\mathcal{S}_{i,CA}^{\uparrow}$ as
$\begin{split}&\mathcal{S}_{i,CA}^{\uparrow}(w)=\begin{cases}(\Sigma_{i}\setminus\rho(E_{H,i}^{\uparrow,a}(w)))\cup(\Sigma_{uc}\cap\Sigma_{i}),&\text{if $w\in\Phi_{i}^{a}(L(H_{i}^{\uparrow})$}\\ \Sigma_{uc}\cap\Sigma_{i},&\text{otherwise }\end{cases}\\ \end{split}$ (15)
@@
Assume that $K$ is conditionally decomposable and Assumption 1 is satisfied. Then the CA-supervisors $\mathcal{S}_{i,CA}^{\uparrow}$ defined in Equation (15) ensure safety of the supervised system under joint sensor and actuator attacks, that is,
@@
(by the definition of $(\mbox{by the definition of }L_{a}(\mathcal{S}_{i,CA}^{\uparrow,a}/G_{i}))$
@@
(by the definition of $(\mbox{by the definition of }\mathcal{S}_{i,CA}^{\uparrow,a}(w))$
@@
(by the definition of $(\mbox{by the definition of }\mathcal{S}_{i,CA}^{\uparrow}(w))$
@@
(because $(\mbox{because }s\in L(H_{i}^{\uparrow})\wedge s\sigma\not\in L(H_{i}^{\uparrow})\Rightarrow\sigma\not\in\Sigma_{uc}\cup\Sigma_{c}^{a})$
@@
(by the definition of $(\mbox{by the definition of }\rho(E_{H,i}^{\uparrow,a}(w)))$
@@
(by letting $(\mbox{by letting }s^{\prime}=s)$
@@ 6 An Illustrative Example
In the figure, states of $G_{a}$ are denoted by $i$, where $i=1,2,3,4,5$, corresponding to the vehicle $V_{a}$ is on Island $i$. Events are $aij$ vehicle $V_{a}$ moves from Island $i$ to Island $j$. Event $aij$ is defined if there is a bridge from Island $i$ to Island $j$. Controllable events are denoted in the figures by small marks on the edges/arrows (that is, ⟶̸⟶$\not\longrightarrow$̸).
@@
$\Sigma=\{aij,bij:\mbox{there is a bridge from Island $i$ to Island $j$}\}.$

### v3

@@ 5 Coordination Control under Cyber Attacks without CA-Controllability and CA-Observability
If $K$ is conditionally decomposable, then for any supervisors $\mathcal{S}_{i}$ for $G_{i}$, $i=1,2$, we have
@@
Replace unobservable transitions in $H_{i}^{\uparrow,a}$ by $\varepsilon$-transitions to obtain
@@
$E_{H,i}^{\uparrow,a}(w)=\{$ $q\in Q_{H,i}^{\uparrow}:(\exists s\in L(H_{i}^{\uparrow}))w\in\Phi_{i}^{a}(s)\wedge\delta_{H,i}^{\uparrow,a}(q_{i,o},s)=q\}.$
Step 4. Construct the CA-observer of $H_{i,\varepsilon}^{\uparrow,a}$ by determinizing $H_{i,\varepsilon}^{\uparrow,a}$ as
Step 5. Design a CA-supervisor $\mathcal{S}_{i,CA}^{\uparrow}$ as
@@
Assume that $K$ is conditionally decomposable. Then the CA-supervisors $\mathcal{S}_{i,CA}^{\uparrow}$ ensure safety of the supervised system under joint sensor and actuator attacks, that is,
@@
$(\mbox{by the definition of }L_{a}(\mathcal{S}_{i,CA}^{\uparrow,a}/G_{i}))$
@@
$(\mbox{by the definition of }\mathcal{S}_{i,CA}^{\uparrow,a}(w))$
@@
$(\mbox{by the definition of }\mathcal{S}_{i,CA}^{\uparrow}(w))$
@@
$(\mbox{because }s\in L(H_{i}^{\uparrow})\wedge s\sigma\not\in L(H_{i}^{\uparrow})\Rightarrow\sigma\not\in\Sigma_{uc}\cup\Sigma_{c}^{a})$
@@
$(\mbox{by the definition of }\rho(E_{H,i}^{\uparrow,a}(w)))$
@@
$(\mbox{by letting }s^{\prime}=s)$
@@ 6 An Illustrative Example
In the figure, states of $G_{a}$ are denoted by $i$, where $i=1,2,3,4,5$, corresponding to the vehicle $V_{a}$ is on Island $i$. Events are aijaij - vehicle $V_{a}$ moves from Island $i$ to Island $j$. Event aijaij is defined if there is a bridge from Island $i$ to Island $j$. Controllable events are denoted in the figures by small marks on the edges/arrows (that is, $\not\longrightarrow$).
@@
[…]